if the point P(x,2) is equidistant from p1(-2,3) and p2( 4,7), find the value of x
Answer by stanbon(75887)

if the point P(x,2) is equidistant from p1(-2,3) and p2( 4,7),
find the value of x
------------
Equation:
(x+2)^2 + (2-3)^2 = (x-4)^2 + (7-2)^2
(x+2)^2 + 1 = (x-4)^2 + 25
----
x^2 + 4x + 5 = x^2 -8x + 41
12x = 36
x = 3
-------------
Cheers,
Stan H.
--------------